Use of methanol has been proposed t rid a wastewater of nitrate by biological denitrification of N2.

a. Write a balanced overall equation for nitrate removal with methanol, using fs(max) and assuming nitrate serves as nutrient source for bacteria.

b. If the (NO3 - ) concentration in the waste water equals 100 mg/L, what concentration of methanol must be added for complete nitrate removal? c. What percentage of the nitrate-nitrogen is used for cell synthesis?
Answer #1

a. the reaction is 5CH3OH + 6NO-> 3N2 + 5CO2 + 7H2O + 6OH-

b. according to the stochiometry of the reaction 5 moles of methanol is needed for 6 moles of nitrate

for each mole of nitrate = 5/6= 0.833 moles of methanol will be needed

for 100mg/L i.e, 0.1g/62=1.61 * 10-3 moles (62 is molar mass of nitrate ion)

1.61 * 10-3 * 0.833= 1.34 * 10-3 moles of methanol

i.e,  1.34 * 10-3 * 32= 42.88 * 10-3g= 42.88mg/L (32 is molar mass of methanol)

42.88mg/L metnaol is required for 100g nitrate